# Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v2

@@ Abstract
The growing number of cases that require digital forensic analysis raises concerns about the ability of law enforcement to conduct investigations promptly. Consequently, this paper delves into the potential and effectiveness of integrating Large Language Models (LLMs) into digital forensic investigation to address these challenges. A comprehensive literature review is carried out, encompassing existing digital forensic models, tools, LLMs, deep learning techniques, and the use of LLMs in investigations. The review identifies current challenges within existing digital forensic processes and explores both the obstacles and possibilities of incorporating LLMs. In conclusion, the study asserts that the adoption of LLMs in digital forensics, with appropriate constraints, has the potential to improve investigation efficiency, improve traceability, and alleviate technical and judicial barriers faced by law enforcement entities.
@@
journal: Elsevier
@@
[label2]organization=Faculty of Law, Criminal Justice and Public Administration, addressline=University of Lausanne, city=Lausanne, country=Switzerland
@@ 1 Introduction
With the pervasive growth of Information and Communication Technology (ICT) and information systems, cybercrimes have seen a significant increase in recent years [1]. As a further compounding factor, the number of “traditional” police investigations that include digital evidence is also increasing [2]. Addressing and investigating this volume of cases presents substantial challenges.
Artificial General Intelligence (AGI) and Large Language Models (LLMs) have become prominent topics of global discussion, prompting researchers to intensify their investigations by leveraging the capabilities of LLMs. The usage of LLMs within the scientific community experienced a rapid surge after 2022, notably with the advent of OpenAI’s ChatGPT platform. In a remarkably short period, this topic has garnered attention from academia, industry, and the research community at large [3]. Simultaneously, researchers are exploring the potential of LLMs across various domains and assessing their impact on the future of science and society. This inquiry also includes an examination of the potential harmfulness associated with the deployment of LLMs [4, 5]. In other words, the use of LLMs in various tasks can be a double-edged sword, necessitating careful consideration depending on the specific situations and contexts.
@@
The main objective of this article is to explore the viability of integrating LLMs into the fundamental phases of the modern DF process. This exploration involved a meticulous and systematic review, adhering to the standards set by Kitchenham [6]. The literature selected for this review was exclusively in English, and non-article sources such as books were excluded. The focus of the review included empirical studies, case studies, and literature reviews within the domain of DF investigation methods, models, frameworks, efficiency, and challenges, as well as LLM capabilities, challenges, and architectures. Additionally, the study underscores the utilisation of autonomous agents based on LLMs, emphasising their capabilities and the potential efficiency gains achievable in the field of DF.
As per the procedure outlined by Kitchenham [6], the quality of the chosen literature was assessed. Design, analysis, and conclusion served as the primary criteria for determining each paper’s quality. In the design benchmark, attention was given to the clarity of the paper’s objectives, definition of variables and measurements, and transparency of methodology. Regarding the analysis benchmark, focus was placed on the description and justification of analytical methods, clarity of the scoring system, and coherence of the analysis. The conclusion benchmark was particularly significant – involving a thorough examination to ensure that the paper’s conclusion adequately addressed study questions, discussed key findings and limitations, and acknowledged any negative aspects.
@@
In light of fast-paced advancements and the recent explosion in LLM-focused research, a substantial influx of LLM-focused research papers has occurred since the launch of ChatGPT in late 2022. Due to this fast pace, many research articles exist solely as preprints on preprint services, e.g., arXiv or OpenReview. To give two examples, the initial papers for GPT-4 [7] and LLaMA [8] are only published on arXiv, but have garnered thousands of citations each. Despite their preprint status, these articles offer essential insights critical for contemporary research and dialogue within the domain, making their incorporation into this article necessary to provide the most up-to-date knowledge and perspectives.
@@
The paper is structured as follows: Section 1.1 provides a comprehensive background for the review, delving into existing DF process models, the challenges inherent in DF, and a detailed overview of the current work conducted with the utilisation of LLMs within DF. In Section 2 the paper delves into the realm of Natural Language Processing (NLP), elucidating the working principles of LLMs, their architectural foundations, and the specifics surrounding specially trained LLMs. Section 3 provides an in-depth review that focuses on the capabilities and benchmark information of LLMs trained for coding tasks, as well as those tailored for vision assistance. Section 3 explores the synergy between DF and LLMs, detailing how LLMs can be effectively employed in each phase of the DF process model. Finally, in Section 6, the paper summarises the future challenges associated with integrating LLMs with automated agents within the DF domain. The conclusion outlines potential avenues for future research and development, shedding light on the path of future DF investigations employing LLMs. The discussion covers not only the potential negative impacts but also the practical difficulties and risks in real world environments.
@@ 1.1 Digital Forensic Context
DF is a process for identifying, preserving, analysing, and documenting digitally recorded data, which originate in electronic devices such as computers, servers, smartphones, and IoT devices [9]. This exercise is required in most criminal cases. Data collected in this process are kept unchanged and safe to present in a court case or to support future investigations conducted by law enforcement agencies [10].
@@ 1.1.1 Digital Forensic Process Models
DF process models consist of a series of activities that help to standardise the investigative process [11] and outline the phases; collection, preservation of evidence, examination or analysis, and reporting.
@@
Casey’s model, as shown in Figure 1, includes phases such as incident recognition, evidence collection, preservation, and presentation, with the examination phase detailed in recovery, harvesting, reduction, and classification [12]. During incident recognition, the focus is on identifying the incident itself, potential evidentiary sources, and expected digital evidence types, as well as delineating the scope of the ensuing investigation. Conversely, investigators systematically acquire pertinent evidence from various sources encompassing computers, smartphones, storage media, and networks during collection and seizure. Preservation assumes paramount importance in upholding the integrity of the evidence, necessitating specific and accurate measures to ensure the unmodified condition of collected data throughout the investigative process. The overarching objective remains the meticulous safeguarding of evidence integrity.
The subsequent phase entails examination, wherein analysts rigorously scrutinise the gathered data to extract pertinent information. This endeavour may involve the employment of diverse forensic hardware and software tools and techniques. The examination process encompasses interpreting extracted information to draw conclusive inferences regarding the events under scrutiny. This phase frequently demands a profound comprehension of both the technology employed and the context surrounding the evidence.
Next is the reporting phase, where the findings derived from the analysis are systematically presented in a format suitable for legal adjudication. This may involve preparing comprehensive reports and providing expert testimony in a court of law. This model emphasises the critical nature of maintaining evidence integrity and provenance, and the requirement for expert analysis to extract and interpret pertinent information. This culminates in a report suitable for legal scrutiny. In particular, the analysis or examination stage is crucial in all models, demanding specialised knowledge in the relevant DF area [17, 18].
@@ 1.1.2 Existing Challenges in Digital Forensics
DF is an evolving field, yet the literature highlights that it still undergoes changes to address ongoing challenges and advancements. Dubey et al. [22] assert that DF faces key challenges, including the complexity of data and its volume, a lack of standardisation, inadequacies in the power of existing tools to support investigations and issues related to timelines.
In addition to the previously mentioned challenges, other issues persist including scope creep in cases due to complexity and the vast data involved, selecting and prioritising the right set of evidence, and efficiently allocating time and investigators for the chosen evidence [23]. Koper et al. [24] focus on a number of these issues from the investigator’s perspective, including challenges in adapting to a system, unexpected time sinks, and frustrations among officers arising from expected operational timelines and the adoption of complex systems. The contemporary landscape of forensic science is characterised by a notable deficiency in proficient agents, exacerbated by the swiftly evolving standards, practices, tools, and techniques within the field. Moreover, the predominant emphasis of law enforcement roles on fieldwork, as opposed to dedicated DF duties, has further contributed to the scarcity of adept human expertise in this domain [25].
@@ 1.1.3 Existing Work With LLMs in Digital Forensics
Scanlon et al. [13] analysed using ChatGPT for DF. In their assessment, the authors evaluated the programming, incident narration, keyword list creation, and DF teaching abilities of ChatGPT. Their conclusion highlighted that while ChatGPT exhibited some hallucinations in the output results, it still serves as an effective assistant for code generation.
Timeline reconstruction helps investigators deduce the chronological “story” of an event. In line with timeline regeneration, Silalahi et al. [28] proposed a method to detect anomalies in a drone flight by employing sentiment analysis with the assistance of a pre-trained LLM. Their approach successfully discerned differences between normal and abnormal events with an accuracy of 92.5%.
Hansken, a DFaaS platform created by the Netherlands Forensic Institute, is designed to help investigators handle evidence and conduct investigations more efficiently [29]. ChatGPT has been used as an assistance for the Hansken DFaaS system, contributing to streamlined processes and improved support for investigators [30]. In a successful experiment, evidence traces were provided to ChatGPT, and it was tasked with analysing the case, ultimately providing conclusions. This utilisation of ChatGPT showcases its potential in assisting with the analytical aspects of investigations, highlighting its ability to process and interpret evidence data [30].
@@ 2 Background of Large Language Models
This section explores LLMs, concentrating on three principal aspects. Initially, it explores the architecture of LLMs, detailing their design and function. Next, it assesses the usability of LLMs, underscoring the features and capabilities that render them apt for a wide range of tasks. Finally, it showcases the versatility of LLMs by discussing their applications across various fields, demonstrating their wide-reaching impact and the extensive scope of their applications.
@@ 2.1 Natural Language Processing
Popular LLMs such as Generative Pre-trained Transformer (GPT) [31], Language Model for Dialogue Applications (LaMDA) [32], Pathways Language Model (PaLM) [33], Bidirectional Encoder Representations from Transformers (BERT) [34], and Large Language Model Meta AI (LLaMA) [8] stem from advances in Natural Language Processing (NLP). NLP, focusing on language-based tasks, utilises both traditional and deep learning models to enable applications such as language translation, text processing, and speech recognition [35].
Deep learning, a branch of machine learning, uses complex computational layers and adaptive weights to improve prediction accuracy, offering a more refined analysis than conventional machine learning [36]. It has excelled in image and speech recognition, and natural language understanding, mimicking the decision-making process of the human brain through artificial neurons. These neurons form networks capable of intricate pattern recognition and data analysis. Central to deep learning are neural networks with multiple hidden layers, which autonomously learn and feature-extract from data, bypassing the need for manual variable selection. This automatic feature extraction makes them exceptionally adept at handling complex tasks [37].
@@ 2.2 LLMs
An LLM is a language model employing neural networks with billions of parameters, trained on extensive text data. These models are engineered to comprehend and generate human language. Fundamentally, they rely on multiple neural network architectures, enabling them to recognise the relationships between words and phrases within sentences [38, 39]. These architectures have been a transformative force in natural language processing. Its capability to excel across a diverse array of language-related tasks distinguishes it as a game-changer, in contrast to being tailored for a singular, specific task.
@@ 2.3.1 Word to Vectors
Word embeddings, as introduced by Mikolov et al. [41], entail precise and high-dimensional vector representations for words, particularly suited for extensive datasets comprising billions of text entries. The authors explored model architectures for word vectorisation, achieving substantial improvements in accuracy while requiring fewer computational resources and reduced training time [41]. In the realm of LLMs, the primary objective is to generate new text based on the extensive dataset, on which it was trained. For this purpose, Vaswani et al. [42] introduced the transformer model, assisted by the word-to-vector model. This architecture incorporates a self-attention mechanism, as well as encoder and decoder processes, enabling the model to rapidly and simultaneously focus on pertinent information.
@@ 2.3.2 Transformer Models
The transformer model initially aimed at machine translation, translating input words into another language, begins with word embedding, where inputs, termed as tokens, are vectorised. Recognising word order is achieved through positional encoding, with two main techniques: absolute and relative. Absolute positional encoding assigns unique vectors to each position, enhancing the model’s ability to recognise word placement and facilitate position-specific attention [43]. Relative positional encoding, on the other hand, calculates the relative positions of words by introducing a bias term that quantifies distances between positions, improving the model’s capacity to understand word relationships within a sequence [43].
Self-attention, a core mechanism within the transformer, calculates the relationship among words in a sentence, allowing the model to assess each word’s similarity to others and generate unique representations for each [44]. The decoder mirrors the encoder’s steps but uses different weights, starting with positional encoding and computing self-attention values to identify the sentence’s initial translation word.
This transformer process, leveraging stacked self-attention and unique positional encoding, has significantly advanced NLP tasks, including machine translation, text generation, and summarisation, by executing these processes in parallel and optimising weights for both encoder and decoder [42, 45].

### v3

@@ Abstract
The ever-increasing workload of digital forensic labs raises concerns about law enforcement’s ability to conduct both cyber-related and non-cyber-related investigations promptly. Consequently, this article explores the potential and usefulness of integrating Large Language Models (LLMs) into digital forensic investigations to address challenges such as bias, explainability, censorship, resource-intensive infrastructure, and ethical and legal considerations.  A comprehensive literature review is carried out, encompassing existing digital forensic models, tools, LLMs, deep learning techniques, and the use of LLMs in investigations. The review identifies current challenges within existing digital forensic processes and explores both the obstacles and the possibilities of incorporating LLMs. In conclusion, the study states that the adoption of LLMs in digital forensics, with appropriate constraints, has the potential to improve investigation efficiency, improve traceability, and alleviate the technical and judicial barriers faced by law enforcement entities.
@@
journal: Forensic Science International: Digital Investigation
@@
[label2]organization=School of Criminal Justice, addressline=University of Lausanne, state=Lausanne, country=Switzerland
[…]
[label3]organization=Present affiliation: Institute of Computer Science, addressline=University of Augsburg, state=Augsburg, country=Germany
@@ 1 Introduction
With the widespread growth of information and communication technology (ICT) and information systems, cybercrimes have seen a significant increase in recent years [1] [^1] https://go.crowdstrike.com/rs/281-OBQ-266/images/GlobalThreatReport2024.pdf. As a further compounding factor, the number of “traditional” police investigations that include digital evidence is also increasing [2]. Addressing and investigating this volume of cases presents substantial challenges.
Generative AI (GenAI) and Large Language Models (LLMs) have become prominent topics of global discussion, prompting researchers to intensify their investigations by leveraging the capabilities of LLMs. The usage of LLMs within the scientific community experienced a rapid surge after 2022, notably with the advent of OpenAI’s ChatGPT platform. In a relatively short period of time, this topic has attracted the attention of academia, industry, and the research community at large [3]. Simultaneously, researchers are exploring the potential of LLMs in various domains and assessing their impact on the future of science and society. This inquiry also includes an examination of the potential harmfulness associated with the deployment of LLMs [4, 5]. In other words, the use of LLMs in various tasks can be a double-edged sword, necessitating careful consideration depending on the specific situations and contexts.
@@
This paper reviews recent advances in the application of LLMs within digital forensics, focussing on established models, methods, and key challenges. By examining contemporary studies from 2019 onwards, the survey highlights core areas, such as automation, investigative methods, and efficiency improvements facilitated by LLMs. In addition, it explores the literature that addresses challenges in both DF and LLMs, covering limitations, ethical considerations, and forensic-specific risks. This comprehensive review synthesises current insights and emerging trends, offering a foundation for understanding the potential and limitations of LLMs in DF contexts.
@@
In light of fast-paced advancements and the recent explosion in LLM-focused research, a substantial influx of LLM-focused research papers has occurred since the launch of ChatGPT in late 2022. Due to this fast pace, many seminal research articles exist solely as preprints on preprint services, e.g., arXiv. To give two examples, the initial papers for GPT-4 [6] and LLaMA [7] are only published on arXiv, but have garnered thousands of citations each. The recent release of DeepSeek-V3 in late 2024 and the corresponding technical reqport [8] is also only available on arXiv. Despite their preprint status, these articles offer essential insights critical for contemporary research and dialogue within the domain, making their incorporation into this article necessary to provide the most up-to-date knowledge and perspectives.
@@
The paper is structured as follows: Section 1.1 provides a comprehensive background for the review, delving into existing DF process models, the challenges inherent in DF, and a detailed overview of the current work conducted with the use of LLMs within DF. In Section 2 the paper delves into the realm of Natural Language Processing (NLP), elucidating the working principles of LLMs, their architectural foundations, and the specifics surrounding specially trained LLMs. Section 3 provides an in-depth review that focusses on the capabilities and benchmark information of LLMs trained for coding tasks, as well as those tailored for vision assistance. Section 3 explores the synergy between DF and LLMs, detailing how LLMs can be effectively employed in each phase of the DF process model. Finally, in Section 6, the paper summarises the future challenges associated with integrating LLMs with automated agents within the DF domain. The conclusion outlines potential avenues for future research and development, shedding light on the path of future DF investigations employing LLMs. The discussion covers not only the potential negative impacts but also the practical difficulties and risks in real world environments.
@@ 1.1 Digital Forensic Context
DF is a process for identifying, preserving, analysing, and documenting digitally recorded data, which originate in electronic devices such as computers, servers, smartphones, and IoT devices [9]. This exercise is required in most criminal cases. Data collected in this process are kept unchanged and safe for presentation in a court case or to support future investigations conducted by law enforcement agencies [10].
@@ 1.1.1 Digital Forensic Process Models
DF process models consist of a series of activities that help standardise the investigative process [11] and outline the phases; collection, preservation of evidence, examination or analysis, and reporting.
@@
Casey’s model, as shown in Figure 1, includes phases such as incident recognition, evidence collection, preservation, and presentation, with the examination phase detailed in recovery, harvesting, reduction, and classification [12]. During incident recognition, the focus is on identifying the incident itself, possible evidentiary sources, and expected digital evidence types, as well as delineating the scope of the ensuing investigation. Conversely, investigators systematically acquire pertinent evidence from various sources encompassing computers, smartphones, storage media, and networks during collection and seizure. Preservation is of paramount importance in upholding the integrity of evidence, necessitating specific and accurate measures to ensure the unmodified condition of the collected data throughout the investigative process. The overarching objective remains the meticulous safeguarding of evidence integrity.
The subsequent phase entails examination, in which analysts rigorously scrutinise the gathered data to extract pertinent information. This endeavour may involve the use of various forensic hardware and software tools and techniques. The examination process includes the interpretation of the information extracted to draw conclusive inferences about the events under scrutiny. This phase often demands a profound understanding of both the technology employed and the context surrounding the evidence.
Next is the reporting phase, where the findings derived from the analysis are presented systematically in a format suitable for legal adjudication. This may involve preparing comprehensive reports and providing expert testimony in a court of law. This model emphasises the critical nature of maintaining the integrity and provenance of the evidence and the requirement for expert analysis to extract and interpret pertinent information. This culminates in a report suitable for legal scrutiny. In particular, the analysis or examination stage is crucial in all models, demanding specialised knowledge in the relevant DF area [17, 18].
@@ 1.1.2 Existing Challenges in Digital Forensics
DF is an evolving field, yet the literature highlights that it still undergoes changes to address ongoing challenges and advancements. Dubey et al. [22] assert that DF faces key challenges, including the complexity of data and its volume, a lack of standardisation, inadequacies in the power of existing tools to support investigations, and issues related to timelines.
In addition to the previously mentioned challenges, other issues persist including scope creep in cases due to complexity and the vast data involved, selecting and prioritising the right set of evidence, and efficiently allocating time and investigators for the chosen evidence [23]. Koper et al. [24] focus on a number of these issues from the investigator’s perspective, including challenges in adapting to a system, unexpected time sinks, and frustrations among officers arising from expected operational timelines and the adoption of complex systems. The contemporary landscape of forensic science is characterised by a shortage of proficient agents, exacerbated by the swiftly evolving standards, practices, tools, and techniques within the field. Moreover, the predominant emphasis of law enforcement roles on fieldwork, as opposed to dedicated DF duties, has further contributed to the scarcity of adept human expertise in this domain [25].
@@ 1.1.3 Existing Work With LLMs in Digital Forensics
Scanlon et al. [13] analysed using ChatGPT for DF. In their assessment, the authors evaluated the programming, incident narration, keyword list creation, and DF teaching abilities of ChatGPT. Their conclusion highlighted that while ChatGPT exhibited some hallucinations in the output results, it still serves as an effective assistant for code generation. Wickramasekara et al. [28] introduced the AutoDFBench benchmarking framework, and corresponding score, to evaluate code generation for DF specific tasks against the tests and datasets used as part of NIST’s Computer Forensics Tool Testing Program (CFTT) [^2] https://www.nist.gov/itl/ssd/software-quality-group/computer-forensics-tool-testing-program-cftt.
Timeline reconstruction helps investigators deduce the chronological “story” of an event. In line with timeline regeneration, Silalahi et al. [29] proposed a method to detect anomalies in a drone flight by employing sentiment analysis with the assistance of a pre-trained LLM. Their approach successfully discerned the differences between normal and abnormal events with an accuracy of 92.5%.
Hansken, a DFaaS platform created by the Netherlands Forensic Institute, is designed to help investigators handle evidence and conduct investigations more efficiently [20]. ChatGPT has been used as an assistance for the Hansken DFaaS system using its bespoke query language, contributing to streamlined processes and improved support for investigators. In these experiments by Henseler and van Beek [30], it was tasked with analysing evidence using Hansken’s trace model. This work demonstrated the potential for ChatGPT in helping with analytical aspects of investigations, highlighting its ability to process and interpret evidence data.
While DF is the main focus of this paper, it is important to recognise the broader application of LLMs in adjacent areas within cybersecurity, many of which overlap with DF. LLMs are proving to be valuable tools in fields such as malware analysis, security log analysis, code security reviews, and intrusion detection areas that bridge the gap between cybersecurity and DF [31, 32, 33]. In malware analysis, LLMs can identify patterns in malicious code, while in log analysis, they assist in detecting anomalies across large datasets, thereby improving response times. In code-related security reviews, LLMs like GRACE have demonstrated the ability to identify vulnerabilities in software, achieving a detection rate of 28.65% of vulnerabilities [34]. These applications contribute to DF investigations by improving and improving evidence collection and analysis.
@@ 2 Background of Large Language Models
This section explores LLMs, concentrating on three principal aspects. Initially, it explores the architecture of LLMs, detailing their design and function. Then it assesses the usability of LLMs, underscoring the features and capabilities that make them suitable for a wide range of tasks. Finally, it showcases the versatility of LLMs by discussing their applications across various fields, demonstrating their wide-reaching impact and the extensive scope of their applications.
@@ 2.1 Natural Language Processing
Popular LLMs such as Generative Pre-trained Transformer (GPT) [35], Language Model for Dialogue Applications (LaMDA) [36], Pathways Language Model (PaLM) [37], Bidirectional Encoder Representations from Transformers (BERT) [38], and Large Language Model Meta AI (LLaMA) [7] stem from advances in Natural Language Processing (NLP). NLP, which focusses on language-based tasks, uses traditional and deep learning models to enable applications such as language translation, text processing, and speech recognition [39].
Deep learning, a branch of machine learning, uses complex computational layers and adaptive weights to improve prediction accuracy, offering a more refined analysis than conventional machine learning [40]. It has excelled in image and speech recognition and natural language understanding, mimicking the decision-making process of the human brain through artificial neurons. These neurons form networks capable of intricate pattern recognition and data analysis. Central to deep learning are neural networks with multiple hidden layers that autonomously learn and extract features from data, bypassing the need for manual variable selection. This automatic feature extraction makes them exceptionally adept at handling complex tasks [41].
@@ 2.2 LLMs
An LLM is a language model that employs neural networks with billions of parameters, trained on extensive text data. These models are engineered to comprehend and generate human language. Fundamentally, they rely on multiple neural network architectures, enabling them to recognise the relationships between words and phrases within sentences [42, 43]. These architectures have been a transformative force in natural language processing. Its capability to excel across a diverse array of language-related tasks distinguishes it as a game-changer, in contrast to being tailored for a singular, specific task.
@@ 2.3.1 Word to Vectors
Word embeddings, as introduced by Mikolov et al. [45], entail precise and high-dimensional vector representations for words, particularly suited for extensive datasets comprising billions of text entries. The authors explored model architectures for word vectorisation, achieving substantial improvements in accuracy while requiring fewer computational resources and reduced training time [45]. In the realm of LLMs, the primary objective is to generate new text based on the extensive dataset on which it was trained. For this purpose, Vaswani et al. [46] introduced the transformer model, assisted by the word-to-vector model. This architecture incorporates a self-attention mechanism, as well as encoder and decoder processes, enabling the model to quickly and simultaneously focus on pertinent information.
@@ 2.3.2 Transformer Models
The transformer model initially aimed at machine translation, translating input words into another language, begins with word embedding, where input, termed tokens, are vectorised. Recognising word order is achieved through positional encoding, with two main techniques: absolute and relative. Absolute positional encoding assigns unique vectors to each position, enhancing the model’s ability to recognise word placement and facilitate position-specific attention [47]. Relative positional encoding, on the other hand, calculates the relative positions of words by introducing a bias term that quantifies the distances between positions, improving the model’s ability to understand the relationships between words within a sequence [47].
Self-attention, a core mechanism within the transformer, calculates the relationship among words in a sentence, allowing the model to assess the similarity of each word with others and generate unique representations for each [48]. The decoder mirrors the encoder’s steps but uses different weights, starting with positional encoding and computing self-attention values to identify the sentence’s initial translation word.
This transformer process, which leverages stacking self-attention and unique positional encoding, has significantly advanced NLP tasks, including machine translation, text generation, and summarisation, by executing these processes in parallel and optimising the weights of both the encoder and decoder [46, 49].